# Mind the Confidence Gap: Overconfidence, Calibration, and Distractor Effects in Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) demonstrate impressive performance across diverse tasks, yet confidence calibration remains a challenge. Miscalibration—where models are overconfident or underconfident—poses risks, particularly in high-stakes applications. This paper presents an empirical study on LLM calibration, examining how model size, distractors, and question types affect confidence alignment. We introduce an evaluation framework to measure overconfidence and investigate whether multiple-choice formats mitigate or worsen miscalibration. Our findings show that while larger models (e.g., GPT-4o) are better calibrated overall, they are more prone to distraction, whereas smaller models benefit more from answer choices but struggle with uncertainty estimation. Unlike prior work, which primarily reports miscalibration trends, we provide actionable insights into failure modes and conditions that worsen overconfidence. These findings highlight the need for calibration-aware interventions and improved uncertainty estimation methods. To ensure reproducibility, we provide our code and dataset [^1] https://github.com/prateekchhikara/llms-calibration.
@@ 1 Introduction
Large Language Models (LLMs) have transformed natural language understanding, achieving state-of-the-art performance across diverse applications, from conversational AI Skjuve et al. (2024); Zhang (2024) to scientific discovery Kumar (2024). Their versatility has also facilitated advancements in multimodal learning, where language models integrate with vision-based systems to enhance tasks such as automated content generation and decision support Zhang et al. (2023); Chhikara et al. (2024). However, as LLMs are increasingly deployed in high-stakes domains such as healthcare, finance, and law, critical concerns arise regarding calibration—the alignment between model confidence and actual correctness Dhuliawala et al. (2023). When LLMs express unwarranted confidence in incorrect predictions, they risk misleading users, spreading misinformation, and reducing trust and reliability in AI-driven systems Chen et al. (2023); Zhang et al. (2024).
Figure 1: LLM generating an incorrect answer with high confidence.
Figure 1 illustrates this issue: when asked “Who received the IEEE Frank Rosenblatt Award in 2010?”, an LLM incorrectly responds with “Geoffrey Hinton”, assigning a 93% confidence score, despite correct answer being “Michio Sugeno”. This overconfidence is especially problematic because users tend to equate high-confidence outputs with reliability Liu et al. (2024). The right-hand side of Figure 1 further demonstrates the broader miscalibration problem: while a well-calibrated model should align confidence with accuracy, real-world LLMs frequently overestimate their correctness.
Although calibration errors in neural networks are well-documented Guo et al. (2017), the calibration behavior of LLMs remains poorly understood, particularly in response to fine-tuning techniques such as Reinforcement Learning from Human Feedback (RLHF) Leng et al. (2024); Li et al. (2024). Prior work Zhu et al. (2023); Geng et al. (2024) has noted that RLHF can amplify overconfidence, but the extent and specific conditions under which this occurs remain largely unexplored. Key open questions include: (i) Do bigger models exhibit better calibration, or does scale introduce new sources of miscalibration? and (ii) Can structured input formats, such as multiple-choice answer choices, improve confidence alignment, or do they introduce additional biases? Addressing these gaps is crucial for ensuring the reliable deployment of LLMs in decision-critical applications.
This paper presents an empirical study of LLM calibration, focusing on three key objectives: (i) Quantifying Overconfidence: We systematically measure miscalibration across multiple LLMs and analyze the extent of overconfidence. (ii) Evaluating the Role of Answer Choices: We assess whether providing structured answer options (e.g., multiple-choice format) improves or worsens accuracy and calibration. (iii) Identifying Failure Modes: We categorize question types where LLMs exhibit the highest miscalibration, revealing systematic weaknesses in uncertainty estimation.
Unlike prior work Ulmer et al. (2024); Kapoor et al. (2024); Jiang et al. (2021) that broadly characterizes miscalibration, our study pinpoints specific conditions that worsen overconfidence, offering an actionable framework for improving LLM reliability. Our findings emphasize the need for task-specific confidence adjustments and structured uncertainty modeling to enhance trust in real-world deployments. Our study reveals that miscalibration is not a uniform problem; its severity varies based on model scale and input structure, with larger models gaining significant calibration improvements when provided with distractors, whereas smaller models struggle with meaningful uncertainty estimation.
@@ 2 Experimental Setup
Evaluation Dataset:
@@
We use the SimpleQA dataset Wei et al. (2024), which provides a reliable benchmark for evaluating LLM factual accuracy and calibration. Comprising short, fact-seeking queries with clearly defined correct answers, SimpleQA enables precise measurement of model confidence and alignment with factual correctness. Its high-quality annotations, verified by multiple independent AI trainers, ensure accuracy and unambiguity, making it well-suited for calibration assessment. The dataset contains 4326 question-answer pairs. For each pair, we used GPT-4o-mini to generate three distractors—factually incorrect yet contextually plausible answers that matched the answer type (e.g., dates for date-based questions), remained distinct, and maintained similar specificity. Additional details on the prompt are provided in Appendix A.
Selected LLMs:
We select models from OpenAI [^2] https://openai.com and GroqCloud [^3] https://groq.com. From OpenAI, we include GPT-4o, GPT-4o-mini, and GPT-4-turbo Hurst et al. (2024), all large-scale models with billions of parameters. From GroqCloud, we select smaller, efficiency-optimized models: LLaMA-3-8B-8192 Dubey et al. (2024), LLaMA3.1-8B-instant, and Gemma2-9B-it Team et al. (2024). More details about the selected models are in Appendix C.
@@
Table 1: Performance metrics of LLMs in the Normal ($\mathcal{N}$) and Distractor ($\mathcal{D}$) settings, including accuracy (correct), NOT_ATTEMPTED (na), ECE, and the number of helped ($\mathcal{D}_{helped}$) and harmed ($\mathcal{D}_{harmed}$) instances.
LLMs $\mathcal{N}_{correct}$ $\mathcal{N}_{na}$ $\mathcal{N}_{ECE}$ $\mathcal{D}_{correct}$ $\mathcal{D}_{na}$ $\mathcal{D}_{ECE}$ $\mathcal{D}_{helped}$ $\mathcal{D}_{harmed}$
@@
GPT-4-turbo [Uncaptioned image] 20.37% 6.17% 0.612 65.40% 0% 0.165 1877 (95.86%) 81 (4.14%)
@@
LLaMA3.1-8b-instant [Uncaptioned image] 5.58% 18.78% 0.799 44.64% 0.12% 0.367 1355 (95.29%) 67 (4.71%)
LLaMA 3-8B-8192 [Uncaptioned image] 4.79% 21.20% 0.810 44.01% 0.55% 0.361 1382 (95.57%) 62 (4.43%)
@@
Figure 2: Reliability diagrams (RDs) showing calibration performance in $\mathcal{N}$ ($\bullet$) and $\mathcal{D}$ ($\bullet$) settings. The numbers on top of bars represent the number of correctly predicted instances. (y-axis: actual accuracy, x-axis: predicted confidence)
Evaluation Criteria
Following prior work, we use GPT-4o-mini as an LLM-based judge to classify responses as CORRECT, INCORRECT, or NOT_ATTEMPTED Packer et al. (2023); Wei et al. (2024). A response is CORRECT if it fully captures the gold target’s key information without contradiction, allowing minor variations in wording, order, or hedging. It is INCORRECT if it contains factual errors, contradictions, or misleading speculation, even if hedged. NOT_ATTEMPTED applies when a response lacks essential information without introducing errors, including vague or evasive answers. We experiment with using the same LLM for both prediction and judgment, finding that smaller LLM judges often misclassify responses or hesitate to assign NOT_ATTEMPTED when no valid answer is generated. Manual inspection confirms these issues, and further details are provided in the Appendix B.
Evaluation Methods:
The standard LLM approach ($\mathcal{N}$) generates answers based only on the given question and a predefined prompt, without external knowledge. We introduce an alternative distractor method ($\mathcal{D}$), where the model receives the question along with the correct answer and three incorrect options (shuffled). This setup enables us to evaluate whether additional context enhances calibration, reinforces overconfidence, or simply encourages random guessing from the provided options. Details on prompts are in the Appendix A.
Evaluation Metrics:

### v2

@@ Abstract
Large Language Models (LLMs) show remarkable proficiency in natural language tasks, yet their frequent overconfidence—misalignment between predicted confidence and true correctness—poses significant risks in critical decision-making applications. We present a comprehensive analysis on calibration in LLMs across nine LLMs and three factual Question-Answering (QA) datasets, systematically comparing standard free-generation settings against structured distractor-augmented prompts. Our evaluation reveals that explicitly incorporating distractors can substantially mitigate miscalibration, achieving relative accuracy improvements up to 460% and ECE reductions up to 90%. Despite general trends, we uncover nuanced findings: large RLHF-tuned models display inherent calibration strengths but can paradoxically suffer increased miscalibration on easier queries, whereas smaller models benefit disproportionately from distractor prompts but remain significantly miscalibrated. Through detailed analyses across question types, we identify persistent calibration failures, particularly in person-based queries. We conclude with concrete recommendations—targeted fine-tuning, structured prompting, and strategic model choice—to ensure reliable, trustworthy LLM deployments. To ensure reproducibility, we provide our code and dataset [^1] https://github.com/prateekchhikara/llms-calibration.
@@
Figure 1: An instance from SimpleQA dataset where an LLM assigns high confidence to an incorrect answer.
@@ 1 Introduction
Large Language Models (LLMs) have significantly advanced natural language understanding, achieving state-of-the-art results across tasks including conversational AI Skjuve et al. (2024); Zhang (2024), scientific discovery Kumar (2024), and multimodal systems Zhang et al. (2023); Chhikara et al. (2024). As LLMs increasingly guide critical decisions in sensitive domains—such as healthcare, finance, and law—the reliability of their confidence estimates becomes paramount. Misalignment between model confidence and actual correctness, known as miscalibration, poses severe risks, potentially eroding user trust and causing costly or hazardous errors Dhuliawala et al. (2023); Geng et al. (2024). For example, as illustrated in Figure 1, when asked “Who received the IEEE Frank Rosenblatt Award in 2010?”, a leading LLM confidently but incorrectly answers “Geoffrey Hinton” with a confidence of 93%, despite the correct answer being “Michio Sugeno”. Such pronounced overconfidence can lead users to mistakenly trust erroneous outputs—particularly problematic in high-stakes applications such as medical diagnoses or financial decisions. Well-calibrated models, on the other hand, report confidence scores accurately reflecting their true reliability, thus enabling systems to flag uncertain predictions for human oversight and significantly mitigating real-world risks.
Modern Question-Answering pipelines adopt “one-right+several-wrong” format — whether through retrieval-augmented candidate spans, knowledge-graph sibling entities, self-consistency checks, or ensemble methods—to improve answer selection. However, these structured, distractor-rich settings introduce novel challenges for confidence estimation—challenges that classical post-hoc calibration methods (temperature scaling, Platt scaling, isotonic regression Guo et al. (2017)) were not originally designed to address. Although such methods have proven effective on small- to medium-scale neural networks, their applicability to today’s large-scale LLMs under real-world, distractor-heavy conditions remains unclear. Prior work has examined isolated factors—model scale, architecture (dense vs. mixture of experts (MoE)), and fine-tuning regime (supervised fine-tuning (SFT) vs. reinforcement learning from human feedback (RLHF)) Leng et al. (2024); Li et al. (2024)—but has not extensively examined how explicit distractors, ubiquitous in deployed QA systems, affect calibration accuracy and confidence ranking for modern LLMs.
Mitigating overconfidence through distractors
Research in cognitive psychology demonstrates that human overconfidence can be reduced by explicitly considering alternative answers before making decisions Lord et al. (1984); Mussweiler et al. (2000). Inspired by this “consider-the-opposite” strategy, we investigate whether presenting LLMs with plausible distractors similarly mitigates their systematic overconfidence and enhances calibration. To be specific, we conduct the first large-scale empirical study of LLM calibration, comparing performance under standard (free-generation) and distractor-augmented settings. Our contributions are fourfold: (1) we introduce a unified calibration benchmark evaluating nine state-of-the-art LLMs—spanning model sizes (8B–70B to greater than 1T), architectures (dense vs. MoE), and fine-tuning methods (SFT vs. RLHF)—across three factual QA datasets (SimpleQA, FaVIQ, TriviaQA); (2) we propose a structured distractor-augmented evaluation paradigm, where models select answers from one correct and multiple plausible incorrect options, enabling simultaneous assessment of accuracy improvements and Expected Calibration Error (ECE); (3) we perform fine-grained analyses across question types (e.g., person, date) to identify conditions of severe miscalibration; and (4) we systematically disentangle how scale, tuning regime, and architecture independently influence model calibration and responsiveness to distractors.
2 Related Work
Intrinsic calibration
Methods directly elicit uncertainty from LLMs. Prompt-based approaches that verbalize confidence Tian et al. (2023); Mielke et al. (2022); Lin et al. or aggregate multiple outputs Xiong et al. have demonstrated effectiveness, particularly for black-box models.
Fine-tuning strategies
Approaches such as calibration-aware RLHF Leng et al. (2024) and Mixup-style data augmentation Park and Caragea (2022) aim to optimize calibration during training. While these methods reduce ECE, they remain sensitive to distribution shifts Liu et al. .
Calibration during pre-training and alignment
Chen et al. (2023) show that calibration emerges early in self-supervised pre-training, and Zhu et al. (2023) demonstrate that instruction tuning and RLHF can preserve or even enhance these gains. Jiang et al. Jiang et al. (2021) find that post-hoc methods like temperature scaling often fail to align confidence with accuracy in factual QA. However, the effects of explicit distractor-based prompting on LLM calibration remain unexplored.
Post-hoc calibration
Techniques adjust predictions after training, with classical approaches like temperature scaling Guo et al. (2017) still underexplored for contemporary LLMs. Recent studies indicate persistent overconfidence even at larger scales, highlighting a potential degradation in calibration performance with increased model size Zhou et al. (2024a, b).
Despite extensive research, existing literature lacks a systematic exploration of structured distractor effects on calibration. Motivated by psychological findings that considering alternative answers can reduce human overconfidence, our work introduces a structured distractor-augmented evaluation framework. Unlike prior methods, we empirically investigate how explicit distractor scenarios—common in practical applications such as retrieval-augmented generation and multiple-choice contexts—impact LLM calibration. Additionally, we conduct detailed, fine-grained analyses across different question types (e.g., person, date, place), uncovering context-dependent calibration challenges previously overlooked. Our findings thus address critical gaps in calibration research, offering practical insights to enhance the reliability and trustworthiness of LLM deployments.
@@ 3 Experimental Setup
3.1 Evaluation Methods
Let our evaluation set be
$\mathcal{S}=\bigl{\{}(q_{i},a_{i})\bigr{\}}_{i=1}^{n},$
where $q_{i}$ is the $i$-th question and $a_{i}$ its ground-truth answer. We compare two prompting regimes:
Free-generation baseline ($\mathcal{N}$)
We prepend each $q_{i}$ with a fixed prompt template $\pi_{N}$ and let model generate a completion answer and the confidence:
$\mathbf{y}_{i}^{(N)},\mathbf{c}_{i}^{(N)}=\mathrm{LLM}\bigl{(}\pi_{N}\,\|\,q_{i}\bigr{)},$
where the model’s final answer is $\mathbf{y}_{i}^{(N)}$ and the associated confidence is $\mathbf{c}_{i}^{(N)}$.
Distractor-augmented setting ($\mathcal{D}$)
For each $(q_{i},a_{i})$ we sample three distractors $\{d_{i,1},d_{i,2},d_{i,3}\}$ and form the choice list
$\mathcal{C}_{i}=\mathrm{shuffle}\bigl{(}\{a_{i}\}\cup\{d_{i,j}\}_{j=1}^{3}\bigr{)}.$
We then feed the model:
$\mathbf{y}_{i}^{(D)},\mathbf{c}_{i}^{(D)}=\mathrm{LLM}\bigl{(}\pi_{D}\,\|\,q_{i}\,\|\,\mathcal{C}_{i}\bigr{)},$
and extract the model’s final answer $\mathbf{y}_{i}^{(D)}$ and record the associated confidence $\mathbf{c}_{i}^{(D)}$. Prompt templates $\pi_{N},\pi_{D}$ are provided in Appendix A.
3.2 Evaluation Datasets:
SimpleQA
@@ 3.2 Evaluation Datasets:
We use the SimpleQA dataset Wei et al. (2024), which provides a reliable benchmark for evaluating LLM factual accuracy and calibration. Comprising short, fact-seeking queries with clearly defined correct answers, SimpleQA enables precise measurement of model confidence and alignment with factual correctness. Its high-quality annotations, verified by multiple independent AI trainers, ensure accuracy and unambiguity, making it well-suited for calibration assessment. The dataset contains 4326 question-answer pairs.
FaVIQ
Park et al. (2022) We select data points from test subset of the R-set. The dataset initially contains 5877 data points, out of which we focus exclusively on the 2922 data points labeled as “supports,” indicating that the provided answer is correct. FAVIQ is particularly appropriate for our experiments due to its construction methodology derived from real-world information-seeking questions. This design inherently reduces strong lexical biases found in other crowdsourced datasets, promoting nuanced semantic understanding.
TriviaQA
We use the TriviaQA dataset Joshi et al. (2017), for evaluating open-domain question answering and factual knowledge retrieval. For our experiments, we select first 1000 question-answer pairs from the validation split of the rc.web.no_content subset, ensuring a diverse yet controlled evaluation set. By restricting our selection to the no-content class, we focus on settings where models must rely purely on prior knowledge without the assistance of retrieved supporting context, isolating intrinsic model calibration behavior.
For each question-answer pair in the above mentioned datasets, we used GPT-4o-mini to generate three distractors—factually incorrect yet contextually plausible answers that matched the answer type (e.g., dates for date-based questions), remained distinct, and maintained similar specificity. Additional details on the prompt are provided in Appendix A.
3.3 Selected LLMs
We select nine representative models, spanning three major families—OpenAI’s GPT-4 series, Meta’s LLaMA lineage, and two leading open-weight models (Gemma-2 and Qwen-qwq). We select these models from OpenAI [^2] https://openai.com and GroqCloud [^3] https://groq.com API services. More details about the selected models and their taxonomy are in Appendix D.
GPT-4 Family
The GPT-4 family consists of three variants: GPT-4o, GPT-4-turbo, and GPT-4o-mini (a smaller, assumed to be 8B-parameter model). All three support an extended 128K-token context window and are instruction-tuned via SFT followed by RLHF to optimize conversational quality. They differ primarily in total parameter count—leading to different trade-offs in inference latency and compute cost.
LLaMA Lineage
Meta’s LLaMA-3 Dubey et al. (2024) series spans three dense checkpoints: an 8B base with 8K window, an 8B “Instant” assistant-tuned model with 128K window, and a 70B base (8K window). Each uses Grouped-Query Attention (GQA) Ainslie et al. for efficient long-context processing; all these variant undergoes SFT and RLHF. In contrast, LLaMA-4-Scout-17b [^4] https://ai.meta.com/blog/llama-4-multimodal-intelligence/ adopts a 16-expert MoE transformer which supports up to 10M tokens, and is fine-tuned with both SFT and RLHF for reasoning.
Open-Weight Alternatives
Gemma2-9b-it Team et al. (2024) leverages knowledge-distillation pre-training on 8T tokens—with interleaved local–global and group-query attention—followed by instruction fine-tuning (SFT + Direct Preference Optimization (DPO) Rafailov et al. (2023)) within 8K-token context. By contrast, Qwen-qwq-32b Hui et al. (2024); Yang et al. (2024) is a 32B-parameter dense model (64 layers, Rotary Positional Embedding (RoPE) Su et al. (2024), SwiGLU Dauphin et al. (2017)) pre-trained on 18T multilingual tokens without RLHF.
3.4 Evaluation Criteria
Following prior work, we use GPT-4o-mini as an LLM-based judge to classify responses as CORRECT, INCORRECT, or NOT_ATTEMPTED Packer et al. (2023); Wei et al. (2024); Chhikara et al. (2025). A response is CORRECT if it fully captures the gold target’s key information without contradiction, allowing minor variations in wording, order, or hedging. It is INCORRECT if it contains factual errors, contradictions, or misleading speculation, even if hedged. NOT_ATTEMPTED applies when a response lacks essential information without introducing errors, including vague or evasive answers. We experiment with using the same LLM for both prediction and judgment, finding that smaller LLM judges often misclassify responses or hesitate to assign NOT_ATTEMPTED when no valid answer is generated. Manual inspection confirms these issues, and further details are provided in the Appendix B.
@@ 3.4 Evaluation Criteria
Table 1: Performance metrics of LLMs in the Normal ($\mathcal{N}$) and Distractor ($\mathcal{D}$) settings on the SimpleQA, FaVIQ, and TriviaQA datasets, including accuracy (correct), NOT_ATTEMPTED (na), ECE, and the number of helped ($\mathcal{D}_{helped}$) and harmed ($\mathcal{D}_{harmed}$) instances with their percentages.
Dataset / LLMs $\mathcal{N}_{correct}$ $\mathcal{N}_{na}$ $\mathcal{N}_{ECE}$ $\mathcal{D}_{correct}$ $\mathcal{D}_{na}$ $\mathcal{D}_{ECE}$ $\mathcal{D}_{helped}$ $\mathcal{D}_{harmed}$
SimpleQA
@@
GPT-4-turbo [Uncaptioned image] 20.37% 6.17% 0.612 65.40% 0.00% 0.165 1877 (95.86%) 81 (4.14%)
@@
LLaMA-3.1-8b-instant [Uncaptioned image] 5.58% 18.78% 0.799 44.64% 0.12% 0.367 1355 (95.29%) 67 (4.71%)
LLaMA-3-8B-8192 [Uncaptioned image] 4.79% 21.20% 0.810 44.01% 0.55% 0.361 1382 (95.57%) 62 (4.43%)
LLaMA-3-70b-8192 [Uncaptioned image] 12.73% 16.46% 0.760 55.81% 0.25% 0.239 1587 (95.09%) 82 (4.91%)
LLaMA-4-scout-17b [Uncaptioned image] 6.70% 8.76% 0.631 50.30% 0.02% 0.285 1763 (95.40%) 85 (4.60%)
@@
Qwen-qwq-32b [Uncaptioned image] 7.59% 3.99% 0.680 51.68% 0.00% 0.253 1784 (96.48%) 65 (3.52%)
FaVIQ
GPT-4o-mini [Uncaptioned image] 47.19% 4.08% 0.426 69.73% 0.24% 0.161 722 (85.85%) 119 (14.15%)
GPT-4-turbo [Uncaptioned image] 54.76% 5.79% 0.357 80.07% 0.31% 0.062 682 (93.81%) 45 (6.19%)
GPT-4o [Uncaptioned image] 56.20% 4.83% 0.315 81.37% 0.21% 0.036 688 (93.73%) 46 (6.27%)
LLaMA-3.1-8b-instant [Uncaptioned image] 36.27% 6.16% 0.532 60.14% 0.31% 0.267 776 (80.83%) 184 (19.17%)
LLaMA-3-8b-8192 [Uncaptioned image] 30.85% 4.81% 0.587 58.53% 0.68% 0.282 892 (86.02%) 145 (13.98%)
LLaMA-3-70b-8192 [Uncaptioned image] 44.95% 4.79% 0.463 72.85% 0.76% 0.139 715 (91.32%) 68 (8.68%)
LLaMA-4-scout-17b [Uncaptioned image] 39.05% 3.74% 0.499 67.85% 0.28% 0.218 861 (89.41%) 102 (10.59%)
Gemma2-9b-it [Uncaptioned image] 35.80% 10.42% 0.581 58.95% 1.28% 0.300 607 (82.25%) 131 (17.75%)
Qwen-qwq-32b [Uncaptioned image] 38.34% 3.62% 0.530 68.54% 0.25% 0.200 850 (92.69%) 67 (7.31%)
TriviaQA
GPT-4o-mini [Uncaptioned image] 81.13% 0.12% 0.104 87.81% 0.00% 0.065 75 (77.32%) 22 (22.68%)
GPT-4-turbo [Uncaptioned image] 90.38% 0.36% 0.025 95.43% 0.00% 0.048 43 (84.31%) 8 (15.69%)
GPT-4o [Uncaptioned image] 89.66% 0.24% 0.071 95.55% 0.00% 0.083 48 (85.71%) 8 (14.29%)
LLaMA-3.1-8b-instant [Uncaptioned image] 75.46% 0.31% 0.153 81.35% 0.00% 0.101 119 (67.61%) 57 (32.29%)
LLaMA-3-8b-8192 [Uncaptioned image] 67.41% 0.51% 0.221 78.93% 0.00% 0.113 166 (74.11%) 58 (25.89%)
LLaMA-3-70b-8192 [Uncaptioned image] 82.86% 0.30% 0.070 91.43% 0.00% 0.026 103 (83.74%) 20 (16.26%)
LLaMA-4-scout-17b [Uncaptioned image] 77.92% 0.40% 0.141 86.87% 0.00% 0.079 127 (74.71%) 43 (25.29%)
Gemma2-9b-it [Uncaptioned image] 70.10% 1.41% 0.230 81.55% 0.00% 0.107 150 (75.76%) 48 (24.24%)
Qwen-qwq-32b [Uncaptioned image] 75.03% 0.62% 0.158 88.08% 0.00% 0.042 133 (86.93%) 20 (13.07%)
@@
Figure 2: Reliability diagrams (RDs) showing calibration performance in $\mathcal{N}$ ($\bullet$) and $\mathcal{D}$ ($\bullet$) settings on the SimpleQA dataset. (y-axis: actual accuracy, x-axis: predicted confidence)
3.5 Evaluation Metrics: